The length of a rectangle is 5 inches more than the width. The perimeter is 34 inches. Find the length and width. Thank you.
Answer by oberobic(2304)

L = length
W = width
L = W + 5 :: given
P = 2L + 2W :: known definition of the perimeter
P = 34 :: given
.
2L + 2W = 34
2(W+5) + 2W = 34
2W + 10 + 2W = 34
4W + 10 = 34
4W = 24
W = 6
.
L = W+5 = 6+5 = 11
.
checking perimeter
2L + 2W = 2(6) + 2(11) = 12 + 22 = 34
OK.
.
Answer:
L = 11
W = 6
.
Done